# Towards Robust Surrogate Models: Benchmarking Machine Learning Approaches to Expediting Phase Field Simulations of Brittle Fracture

### v1

@@ 3.4 Ensemble approaches
All ensemble results shown in Fig. 9b and Fig. 10b and listed in Table 2 reflect these threshold combinations. As widely observed in the literature [5, 82], ensembling improves the prediction performance of both model types. However, the magnitude of this improvement differs between the two. For the FNO predictions, Hard voting improved the performance of the model by $\approx 10\%$ to 0.6831, the soft voting by $\approx 6.8\%$ to 0.6804 and the stacking achieved the highest boost, improving the performance by $\approx 18.1\%$ to 0.7333. The stacking method with non linearity through ReLU activation function and context aware filters as described in Section 2.4 outperformed the other ensembling methods. One possible explanation for this outcome is that training a meta learner like the stacker model allows the network to act as a localized router learning which base model’s output to prioritize for different types of crack patterns. For example, one initial model might perform better on horizontal cracks while another may more accurately capture crack coalescence. In the future, we could also integrate this stacking model into the architecture itself to form a simple Mixture of Experts (MoE) model [109]. Looking at both Figs. 9 and 10, we can see that the contextual filters of the stacking model helped to filter out False positives and thus considerably improved the prediction accuracy.
@@ 3.5 Overall Baseline Model Performance – Methods Comparison
When comparing the performance of the baseline FNOs to the baseline UNet, we can see that the FNOs produced more smoothed out predictions and struggled to accurately model high frequency features of the domain such as the crack paths. In contrast, the UNet models with Focal-Dice loss function, generally achieved higher Dice scores on the dataset compared to the FNO models and produced sharper predictions. The ensembling methods resulted in a bigger relative improvement on the FNO predictions of approximately $18\%$ when using the stacking method achieving a comparable performance to the UNet models with a Dice score of 0.7333. This highlights the potential of ensembling methods to mitigate some of the limitations of operator learning architectures especially in challenging prediction tasks.
That being said, there are several promising directions that could be explored to improve upon the baseline model performance and reliability from both the FNO and UNet starting points presented here. First, adopting more advanced UNet-based architectures such as ConvNet UNets [110] or Multires-WNet [111], which extract multi-scale features, could lead to higher performance on these data. Another promising avenue is to incorporate more robust autoregressive training strategies, such as push forward and temporal bundling techniques introduced by Brandstetter et al. [112], to stabilize training over time sequences. Lastly, increasing the number of retained Fourier modes in the FNO architecture may further improve its capacity to model high frequency features and help improve baseline performance. Overall, we anticipate significant future work in this direction. The purpose of this paper is to showcase a challenging problem, and provide a starting point upon which future improvements can be made.
@@ 3.6 A Note on Performance Metrics
In this study, we have addressed the limitation of MSE by using the Sørensen–Dice score for evaluating the performance on the phase field channel of the predictions – in the literature also known as the $F_{1}$ score. This metric is used very commonly in semantic segmentation and classification tasks. This score can be calculated as follows:
$DSC=\frac{2|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|}=\frac{TP}{TP+\frac{1}{2}(TP+FP)}$ (23)
where the $\hat{Y}$ is the binarized prediction field and $Y$ is the binarized target field. Here we can also see the metric in the context of binary classification where TP stands for True Positive and FP stands for False Positive. The main motivation behind choosing this metric is that the crack patterns occupy a very small portion of the field (approximately $2\%$ of the pixels), with the majority of pixels belonging to the background. This creates a highly imbalanced dataset, making this metric well suited for evaluating performance in fields with sparse, intricate patterns where the true negative classes, the undamaged regions, occupy a large portion of the domain. As can be seen in equation 23, the Dice score is not affected by the true negatives. While MSE might be a suitable choice for smooth, dense and continuous by nature fields such as displacement fields providing a reliable measure of overall field accuracy, it tends to be dominated by undamaged or background regions of the domain in the phase field channel of the data, making it insensitive to small features of the domain.
While Dice score provides a useful measure for this case, it has notable limitations. It does not penalize the discontinuous crack paths, for example in Fig. 9 b all of the samples other than the best performing one shows visible discontinuities in the crack trajectory, and also can produce higher scores for samples with more initial cracks or those where the cracks travel a longer and more complex trajectory within the domain. This behavior can be seen in Fig. 9b when comparing the best and worst examples. Samples that have more initial cracks tend to get a higher score in comparison to those with a smaller number of initial cracks. These limitations highlight the need for more physically grounded or geometry aware evaluation metrics that account for topological correctness, connectivity, and adherence to physics based constraints. For example, future studies could explore alternative hybrid metrics that combine pixel overlap with physics informed loss components to better assess prediction quality in fracture simulations.
@@ 4 Conclusion
Despite achieving reasonable results with two of the baseline models, we observed that these models are highly sensitive to initialization and are prone to converging to different local minima. This behavior reflects the inherent non convexity of the underlying optimization problem and suggests that our dataset can serve as a challenging benchmark for evaluating the robustness of different modeling approaches. In particular, it provides a valuable opportunity to study the effects of training hyperparameters, random seeds, loss functions and training strategies on model performance and convergence. We can also anticipate future studies in several directions, including model calibration, development of more effective ensembling strategies, design of better metrics and loss functions that can help with spectral bias in neural networks by helping to capture high frequency features such as crack paths and the effects of data augmentations based on generative AI models.

### v2

@@ 3.4 Ensemble approaches
All ensemble results shown in Fig. 10b and Fig. 11b and listed in Table 5 reflect these threshold combinations. As widely observed in the literature [8, 92], ensembling improves the prediction performance of both model types. However, the magnitude of this improvement differs between the two. For the FNO predictions, Hard voting improved the performance of the model by $\approx 10\%$ to 0.6831, the soft voting by $\approx 9.6\%$ to 0.6804 and the stacking achieved the highest boost, improving the performance by $\approx 18.1\%$ to 0.7333. The stacking method with non-linearity through ReLU activation function and context aware filters as described in Section 2.4 outperformed the other ensembling methods. One possible explanation for this outcome is that training a meta learner like the stacker model allows the network to act as a localized router learning which base model’s output to prioritize for different types of crack patterns. For example, one initial model might perform better on horizontal cracks while another may more accurately capture crack coalescence. In the future, we could also integrate this stacking model into the architecture itself to form a simple Mixture of Experts (MoE) model [120]. Looking at both Figs. 10 and 11, we can see that the contextual filters of the stacking model helped to filter out False positives and thus considerably improved the prediction accuracy.
@@ 3.5 Overall Baseline Model Performance – Methods Comparison
When comparing the performance of the baseline FNOs to the baseline UNet, we can see that the FNOs produced more smoothed out predictions and struggled to accurately model high frequency features of the domain such as the crack paths. This spectral bias is a well known limitation of the FNO architecture [115]. However, in our experiments, the models were still able to capture certain high frequency features in both the phase field and displacement fields as illustrated in the top row of Fig. 14. In contrast, the UNet models with Focal-Dice loss function, generally achieved higher Dice scores on the dataset compared to the FNO models and produced sharper predictions. The Dice score achieved by the UNet models without ensembling (0.6804) shows a moderately good overlap between the predicted crack patterns and the ground truth. Given the sparsity of the cracked regions, this score represents a non-trivial result and confirms that the UNet models are learning a representation of the underlying pattern, though considerable room for improvement remains. The ensembling methods resulted in a larger relative improvement on the FNO predictions of approximately $18\%$ when using the stacking method achieving a comparable performance to the UNet models with a Dice score of 0.7333. This highlights the potential of ensembling methods to mitigate some of the limitations of operator learning architectures especially in challenging prediction tasks.
That being said, there are several promising directions that could be explored to improve upon the baseline model performance and reliability from both the FNO and UNet starting points presented here. First, adopting more advanced UNet-based architectures such as ConvNet UNets [121] or Multires-WNet [122], which extract multi-scale features, could lead to higher performance on these data. Another promising avenue is to incorporate more robust autoregressive training strategies, such as push forward and temporal bundling techniques introduced by Brandstetter et al. [123], to stabilize training over time sequences. Lastly, increasing the number of retained Fourier modes in the FNO architecture may further improve its capacity to model high frequency features and help improve baseline performance. The purpose of this paper is to showcase a challenging problem, and provide a starting point upon which future improvements can be made.
@@ 3.6 A Note on Performance Metrics
The limitations of the MSE are particularly pronounced in the phase field channel of the dataset, where the number of cracked pixels is very small compared to intact regions, resulting in highly imbalanced data. To better evaluate the performance of the crack pattern prediction, we use the Sørensen–Dice score, also known as the $F_{1}$ score in the literature. We note that the Dice score cannot be used effectively on the continuous values such as those found in the displacement field, but can be applied to the binary classification of each pixel in the predicted phase field channel as either damaged or undamaged. This metric is used very commonly in semantic segmentation and classification tasks. This score can be calculated as follows:
$DSC=\frac{2|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|}=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$ (23)
where the $\hat{Y}$ is the binarized prediction field and $Y$ is the binarized target field. Here we can also see the metric in the context of binary classification where TP stands for True Positive, FP stands for False Positive and FN stands for False Negative. The main motivation behind choosing this metric is that the crack patterns occupy a very small portion of the field (approximately $2\%$ of the pixels), with the majority of pixels belonging to the background. This creates a highly imbalanced dataset, making this metric well suited for evaluating performance in fields with sparse, intricate patterns where the true negative classes, the undamaged regions, occupy a large portion of the domain. As can be seen in equation 23, the Dice score is not affected by the true negatives. While MSE might be a suitable choice for smooth, dense and continuous by nature fields such as displacement fields providing a reliable measure of overall field accuracy, it tends to be dominated by undamaged or background regions of the domain in the phase field channel of the data, making it insensitive to small features of the domain.
While Dice score provides a useful measure for this case, it has notable limitations. It does not penalize the discontinuous crack paths, for example in Fig. 10 b all of the samples other than the best performing one show visible discontinuities in the crack trajectory, and also can produce higher scores for samples with more initial cracks or those where the cracks travel a longer and more complex trajectory within the domain. This behavior can be seen in Fig. 10b when comparing the best and worst examples. Samples that have more initial cracks tend to get a higher score in comparison to those with a smaller number of initial cracks. These limitations highlight the need for more physically grounded or geometry aware evaluation metrics that account for topological correctness, connectivity, and adherence to physics based constraints. For example, future studies could explore alternative hybrid metrics that combine pixel overlap with physics informed loss components to better assess prediction quality in fracture simulations.
@@ 4 Conclusion
Despite achieving reasonable results with two of the baseline models, we observed that these models are highly sensitive to initialization and are prone to converging to different local minima. This behavior reflects the inherent non-convexity of the underlying optimization problem and suggests that our dataset can serve as a challenging benchmark for evaluating the robustness of different modeling approaches. In particular, it provides a valuable opportunity to study the effects of training hyperparameters, random seeds, loss functions and training strategies on model performance and convergence. We anticipate several promising directions for future research, including the design of more data efficient models through tighter integration of governing equations [23], developing foundational models for physics [127, 128, 129] and creating more effective ensembling strategies [115]. Further work could involve designing better metrics and loss functions to mitigate spectral bias[130], and training strategies to achieve a more stable rollout for autoregressive models [123, 131, 132]. We also encourage the expansion of the dataset to more complex 2D and 3D fracture scenarios and have open-sourced our code to facilitate this effort.
@@ Appendix B Additional results of evaluating FNO model on the test dataset
In Fig. 15-17, we show the ground truth, the prediction and the absolute difference of the prediction of each channel from the ground truth for each FNO seed.
Figure 15: This figure presents the initial condition, ground truth and predicted phase field and the absolute difference of the prediction of the phase field channel from the ground truth for each FNO seed.
Figure 16: This figure presents the ground truth and the predicted X component of the displacement field and the absolute difference of the prediction of the X component of the displacement field channel from the ground truth for each FNO seed.
Figure 17: This figure presents the ground truth and the predicted Y component of the displacement field and the absolute difference of the prediction of the Y component of the displacement field channel from the ground truth for each FNO seed.